# A New Fast Finite Difference Scheme for Tempered Time Fractional Advection-Dispersion Equation with a Weak Singularity at Initial Time

### v2

@@ 3.2. Stability of difference schemes
$+\frac{1}{\Gamma(1-\alpha)}\left[a_{0,n}\hat{\rho}^{n}[k]+\sum\limits_{l=1}^{n-1}\left(a_{n-l,n}+b_{n-1-l,n}\right)\hat{\rho}^{l}[k]+(b_{n-1,n}+\frac{e^{-\lambda t_{n+\frac{1}{2}}}}{t_{n+\frac{1}{2}}^{\alpha}})\hat{\rho}^{0}[k]\right]$ (3.9)
@@
$\left|\left[\frac{1}{\tau_{n+1}}-\frac{1-2e^{-\lambda\frac{\tau_{n+1}}{2}}}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}-\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{\Gamma(1-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}+a_{0,n}\right]\hat{\rho}^{n}[k]+\right$.
$\left.\frac{1}{\Gamma(1-\alpha)}\sum\limits_{l=1}^{n-1}\left(a_{n-l,n}+b_{n-1-l,n}\right)\hat{\rho}^{l}[k]+\frac{1}{\Gamma(1-\alpha)}(b_{n-1,n}+\frac{e^{-\lambda t_{n+\frac{1}{2}}}}{t_{n+\frac{1}{2}}^{\alpha}})\hat{\rho}^{0}[k]\right|$
$\leq\left[\frac{1}{\tau_{n+1}}-\frac{1-2e^{-\lambda\frac{\tau_{n+1}}{2}}}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}-\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{\Gamma(1-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right]|\hat{\rho}^{0}[k]|+$
@@
Using the triangular inequality
$\frac{\left|-\frac{1-cos(h\beta)}{h^{2}}-\frac{1}{2h}isin(h\beta)+\left(\frac{1}{\tau_{n+1}}-\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)\right|+\frac{1}{\Gamma(2-\alpha)}\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{(\frac{\tau_{n+1}}{2})^{\alpha}}}{\left|\frac{1-cos(h\beta)}{h^{2}}+\frac{1}{2h}isin(h\beta)+\left(\frac{1}{\tau_{n+1}}+\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)\right|}$
$\leq\frac{\left|-\frac{1-cos(h\beta)}{h^{2}}+\left(\frac{1}{\tau_{n+1}}-\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)\right|+\frac{1}{2h}|sin(h\beta)|+\frac{1}{\Gamma(2-\alpha)}\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{(\frac{\tau_{n+1}}{2})^{\alpha}}}{\left|\frac{1-cos(h\beta)}{h^{2}}+\frac{1}{2h}isin(h\beta)+\left(\frac{1}{\tau_{n+1}}+\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)\right|}$
When $\frac{1-cos(h\beta)}{h^{2}}+\left(\frac{1}{\tau_{n+1}}-\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)\geq 0$,
$\frac{\left|-\frac{1-cos(h\beta)}{h^{2}}+\left(\frac{1}{\tau_{n+1}}-\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)\right|+\frac{1}{2h}|sin(h\beta)|+\frac{1}{\Gamma(2-\alpha)}\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{(\frac{\tau_{n+1}}{2})^{\alpha}}}{\left|\frac{1-cos(h\beta)}{h^{2}}+\frac{1}{2h}isin(h\beta)+\left(\frac{1}{\tau_{n+1}}+\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)\right|}$
$\leq\frac{-\frac{1-cos(h\beta)}{h^{2}}+\left(\frac{1}{\tau_{n+1}}-\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)+\frac{1}{2h}|sin(h\beta)|+\frac{1}{\Gamma(2-\alpha)}\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{(\frac{\tau_{n+1}}{2})^{\alpha}}}{\frac{1-cos(h\beta)}{h^{2}}+\left(\frac{1}{\tau_{n+1}}+\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)}$
$\leq 1$
$\iff\frac{1}{2h}|sin(h\beta)|+\frac{1}{\Gamma(2-\alpha)}\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{(\frac{\tau_{n+1}}{2})^{\alpha}}\leq 2\frac{1-cos(h\beta)}{h^{2}}+\frac{1}{\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}$
Due to our assumption, we have $2\frac{1-cos(h\beta)}{h^{2}}+\frac{1}{\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\geq\frac{2}{\tau_{n+1}}$. To prove our conclusion, we only need
$\frac{1}{2h}|sin(h\beta)|+\frac{1}{\Gamma(2-\alpha)}\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{(\frac{\tau_{n+1}}{2})^{\alpha}}\leq\frac{2}{\tau_{n+1}}$ (3.13)
When $\frac{1-cos(h\beta)}{h^{2}}+\left(\frac{1}{\tau_{n+1}}-\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)\leq 0$,
$\frac{\left|-\frac{1-cos(h\beta)}{h^{2}}+\left(\frac{1}{\tau_{n+1}}-\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)\right|+\frac{1}{2h}|sin(h\beta)|+\frac{1}{\Gamma(2-\alpha)}\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{(\frac{\tau_{n+1}}{2})^{\alpha}}}{\left|\frac{1-cos(h\beta)}{h^{2}}+\frac{1}{2h}isin(h\beta)+\left(\frac{1}{\tau_{n+1}}+\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)\right|}$
$\leq\frac{\frac{1-cos(h\beta)}{h^{2}}-\left(\frac{1}{\tau_{n+1}}-\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)+\frac{1}{2h}|sin(h\beta)|+\frac{1}{\Gamma(2-\alpha)}\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{(\frac{\tau_{n+1}}{2})^{\alpha}}}{\frac{1-cos(h\beta)}{h^{2}}+\left(\frac{1}{\tau_{n+1}}+\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right)}$
$\leq 1$
$\iff$ $\frac{1}{2h}|sin(h\beta)|+\frac{1}{\Gamma(2-\alpha)}\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{(\frac{\tau_{n+1}}{2})^{\alpha}}\leq\frac{2}{\tau_{n+1}}$ (3.14)
Then, under the assumption of equation (3.13)(3.14), we get $|\hat{\rho}^{n+1}[k]|\leq|\hat{\rho}^{0}[k]|$ by equation (3.9)(3.11) and (3.12). By using mathematical induction, we complete the proof. ∎

### v3

@@ 3.2. Stability of difference schemes
$+\frac{1}{\Gamma(1-\alpha)}\left[a_{0,n}\hat{\rho}^{n}[k]+\sum\limits_{l=1}^{n-1}\left(a_{n-l,n}+b_{n-1-l,n}\right)\hat{\rho}^{l}[k]+(b_{n-1,n}+\frac{e^{-\lambda t_{n+\frac{1}{2}}}}{t_{n+\frac{1}{2}}^{\alpha}})\hat{\rho}^{0}[k]\right]$
$=\left[\frac{1}{h^{2}}cos(h\beta)-\frac{1}{2h}isin(h\beta)+\left(\frac{1}{\tau_{n+1}}-\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}-\frac{1}{h^{2}}\right)\right]\hat{\rho}^{n}[k]$
$+\frac{1}{\Gamma(1-\alpha)}\left[(a_{0,n}+\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{(1-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}-\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{(\frac{\tau_{n+1}}{2})^{\alpha}})\hat{\rho}^{n}[k]\right$.
$+\left.\sum\limits_{l=1}^{n-1}\left(a_{n-l,n}+b_{n-1-l,n}\right)\hat{\rho}^{l}[k]+(b_{n-1,n}+\frac{e^{-\lambda t_{n+\frac{1}{2}}}}{t_{n+\frac{1}{2}}^{\alpha}})\hat{\rho}^{0}[k]\right]$ (3.9)
@@
[…]
$+\left.\left.\sum\limits_{l=1}^{n-1}\left(a_{n-l,n}+b_{n-1-l,n}\right)\hat{\rho}^{l}[k]+(b_{n-1,n}+\frac{e^{-\lambda t_{n+\frac{1}{2}}}}{t_{n+\frac{1}{2}}^{\alpha}})\hat{\rho}^{0}[k]\right]\right|$
$\leq\left[\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}-\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{\Gamma(1-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\right]|\hat{\rho}^{0}[k]|+$
@@
Combine 3.9,3.11 and 3.12, we have
$\left|-\frac{1}{h^{2}}cos(h\beta)+\frac{1}{2h}isin(h\beta)+\left(\frac{1}{\tau_{n+1}}+\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}+\frac{1}{h^{2}}\right)\right|\left|\hat{\rho}^{n+1}[k]\right|$
$\leq$ $\left|\frac{1}{h^{2}}cos(h\beta)-\frac{1}{2h}isin(h\beta)+\left(\frac{1}{\tau_{n+1}}-\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}-\frac{1}{h^{2}}\right)\right|\left|\hat{\rho}^{n}[k]\right|$
$+\frac{e^{-\lambda\frac{\tau_{n+1}}{2}}}{\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}}\left|\hat{\rho}^{0}[k]\right|$ (3.13)
Let
$A$ $=\frac{1-cos(h\beta)}{h^{2}},B=\frac{1}{2\Gamma(2-\alpha)(\frac{\tau_{n+1}}{2})^{\alpha}},C=2Be^{-\lambda\frac{\tau_{n+1}}{2}};$
$a$ $=-A+\frac{1}{\tau_{n+1}}-B,b=A+\frac{1}{\tau_{n+1}}+B,d=\frac{sin(h\beta)}{2h},$
then we want to proof the following inequality
$\sqrt{a^{2}+d^{2}}+C\leq\sqrt{b^{2}+d^{2}}$
$\iff$ $\sqrt{a^{2}+d^{2}}+\sqrt{b^{2}+d^{2}}\leq\frac{(b-a)(b+a)}{C}$
$\iff$ $a^{2}+d^{2}\leq\frac{(b-a)^{2}(b+a)^{2}}{C^{2}}-\frac{2(b-a)(b+a)}{C}\sqrt{b^{2}+d^{2}}+b^{2}+d^{2}$
$\iff$ $\sqrt{b^{2}+d^{2}}\leq\frac{(b-a)(b+a)}{2C}+\frac{C}{2}$
$\iff$ $(A+B)^{2}+\frac{1}{\tau_{n+1}^{2}}+d^{2}\leq B^{2}e^{-\lambda\tau_{n+1}}+\frac{(A+B)^{2}}{B^{2}e^{-\lambda\tau_{n+1}}\tau_{n+1}^{2}}$
$\iff$ $d^{2}\leq\frac{1-B^{2}e^{-\lambda\tau_{n+1}}\tau_{n+1}^{2}}{B^{2}e^{-\lambda\tau_{n+1}}\tau_{n+1}^{2}}((A+B)^{2}-B^{2}e^{-\lambda\tau_{n+1}})$
noticing that $d^{2}=\frac{sin(h\beta)^{2}}{4h^{2}}\leq\frac{1-cos(h\beta)^{2}}{4h^{2}}\leq\frac{A}{2}$ and $(A+B)^{2}-B^{2}e^{-\lambda\tau_{n+1}}\geq A^{2}+2AB$, so $\frac{(A+B)^{2}-B^{2}e^{-\lambda\tau_{n+1}}}{d^{2}}\geq 2\frac{A^{2}+2AB}{A}\geq 4B$ which means we only need
$1\leq 4B\frac{1-B^{2}e^{-\lambda\tau_{n+1}}\tau_{n+1}^{2}}{B^{2}e^{-\lambda\tau_{n+1}}\tau_{n+1}^{2}}$ (3.14)
Combining the obtained expression with the condition $\frac{1}{2}<\Gamma(2-\alpha)<1$, we employ bounding arguments to simplify the inequality 3.14, and it is therefore sufficient to focus on the relation $(\frac{\tau_{n+1}}{2})^{2-2\alpha}<\frac{1}{3}$ which is roughly estimated and easily ensured in numerical simulation. Finally, we get $|\hat{\rho}^{n+1}[k]|\leq|\hat{\rho}^{0}[k]|$ by equation (3.13) and (3.14). By using mathematical induction, we complete the proof. ∎